# Physics-informed machine learning for combustion: A review

### v1

@@ 4.2 Surrogate models of chemical kinetics
In combustion simulations, the direct integration of the chemical reaction source terms is computationally expensive and is the main bottleneck for fast simulations of industrial combustion. Therefore, many methods have been proposed to accelerate the computation 43, 44, 113, 269, such as tabulation methods, reduced chemistry, and manifold methods. ML methods have also been adopted for computation acceleration, usually by constructing NN surrogate models of the chemical kinetics 43. The NNs are usually first trained using kinetic data, and then used to replace the integration of reaction source terms, so the overall computation efficiency can be improved. Based on the type of NN outputs, the NN surrogate models can be classified into two paradigms: discrete and continuous models. Specifically, the discrete models learn the mapping between the state variables at the current time and next time (i.e., following the AR paradigm), or the corresponding changes:
@@
Almeldein et al. 267 developed the physics-informed surrogate NNs for accelerating chemical kinetics calculations following Equation (4.12), where the conservation of four elements was added in the loss function. The mixture-of-experts architecture was adopted to allow multiple subnetworks to specialize in different regimes, such as pre-ignition low-temperature chemistry, high-temperature reactions, and post-combustion equilibriums. The 0D CH4-air combustion was tested, demonstrating the effectiveness of element conservation and the acceleration ability of the surrogate models on both CPUs and GPUs. Zhang et al. 268, 285 proposed the parametric physics-informed surrogate NNs for combustion reaction kinetics (CRK-PINNs) based on the paradigm of Equation (4.12). As illustrated in Figure 8, the kinetic ODEs and the conservations of elements, species, and energy were considered in the loss function to improve the physical completeness and low data dependence. A logarithmic normalization strategy was adopted to mitigate the multiscale of species and stiffness of the system. Based on the H2-air mechanism with 10 species and 21 reactions, the CRK-PINN was validated in three scenarios: solving forward and inverse problems, surrogate chemical kinetics, and accelerating simulations of combustion reacting flows. In the first scenario, CRK-PINN can solve the ignition process without data, reconstruct the process with sparse data, and infer intermediate species and temperature with the data of major species. In the second scenario, CRK-PINN can surrogate chemical kinetics in wide operating conditions and exhibited higher accuracy and lower data dependence than data-driven NNs for both time-continuous and time-advanced predictions. In the last scenario, CRK-PINN can replace the direct integration solver to predict the reaction source terms in the DNS of a 2D laminar Bunsen flame and a 3D turbulent jet flame, showing superior accuracy than data-driven NNs and great acceleration compared to the direct integration method.
Figure 8: CRK-PINN architecture. Reprinted from Reference 268, with permission.
In addition to the soft constraints in the form of loss functions, there are also studies that impose hard constraints on physical conservations. Zhang et al. 114 proposed a multiscale sampling method to enhance the accuracy and robustness of NNs to surrogate chemical kinetics, showing superior ability than the manifold, Monte Carlo, and GAN sampling. The Box-Cox transformation (BCT) was adopted to preprocess the data of species mass fractions:
@@
which is a good alternative to the logarithmic normalization 268. Although the paradigm of Equation (4.13) was adopted, only the species were directly predicted by the NNs, while the temperature and pressure were updated by transport fluxes, thus ensuring the mass and energy conservation. The algorithm has been validated on various CFD cases of combustion and been applied to DeepFlame 115, a deep learning empowered open-source platform for reacting flow simulations. Wang et al. 113 extended the method by imposing constraints on element conservations in both soft and hard ways. As illustrated in Figure 9, the ANN-soft penalized the element conservations by adding a loss term, while ANN-hard enforced the element conservations by adding a correction term, $\Delta\mathbf{Y}_{c}$, to the NN-predicted mass fractions. The correction term can be easily derived by the element conservations and the changes of species mass fractions, and was first computed in the Box-Cox transformed space to avoid the multiscale issue. One advantage of ANN-hard over ANN-soft is that there is no cumbersome adjustment of the loss weights. In terms of prediction accuracy and error of element conservations, ANN-hard exhibited superior performance than ANN-soft and they all outperformed the regular ANN on various flame cases. The computational overhead caused by the correction term was negligible, demonstrating the efficiency of ANN-hard. In fact, the correction method to ensure element conservations has been adopted in earlier studies 51, 52, where the NNs were trained to surrogate the kinetics of reduced chemistry, following the paradigm of Equation (4.15).s
@@
Figure 9: ANN-hard: schematic and comparisons with other approaches. Reprinted from Reference 113, with permission.
@@
In addition to the paradigm of function learning, some studies follow the paradigm of operator learning to learn combustion chemical kinetics, using neural operators such as the DeepONet 95 and FNO 96. Some data-driven DeepONets have been used to surrogate chemical kinetics 286, 139, 140, where the input of branch nets and the final output are similar to Equation (4.12), while the input of truck nets is related to the temporal coordinates or increments, enabling the models to learn kinetics with variable $\Delta t$, which is often fixed in classical surrogate models. To some extent, they can be viewed as a hybrid paradigm of AR and NAR learning. One insight that these works can provide is that learning kinetics in the latent space with reduced dimension can improve computational efficiency, which can be achieved by integrating a self-supervised autoencoder (AE) 139, 140, as shown in Figure 10. Kumar et al. 140 also accelerated the computation by selecting representative species for kinetics learning. The other scalars can be reconstructed by a pre-trained correlation NN. This strategy was also adopted in Reference 102, 103, where the species and element conservations were integrated into the loss functions of both correlation NN and DeepONet, thus constituting a physics-informed learning. The framework was tested for CH4 oxidation, demonstrating that the physical constraints can reduce the conservation errors without loss of prediction accuracy. Zanardi et al. 104 used the PI-DeepONets to learn the solution operator of non-equilibrium chemistry in hypersonic flows, where the coarse-grained (CG) kinetic ODEs were considered. The input of the branch and truck nets were the ICs and temporal coordinates, respectively. The framework was tested on the O2–O mixture, exhibiting two orders of magnitude in speedup compared to numerical integrators. Later, Zanardi et al. 105 extended the framework to be more physics-informed by many strategies, including a Boltzmann transformation layer to enforce local equilibrium distributions between states in the same group, the hierarchical architecture and transfer learning between different temporal scales, and the adaptive online pruning technique. Weng et al. 269 applied FNO to learn stiff chemical kinetics, following the paradigm of Equation (4.12). The loss of element and species conservations were considered and the BCT (Equation (4.16)) was also adopted. Based on the focal loss 287, a balanced loss function was proposed to address the unbalanced sampling issue in complex reacting flow problems. The results on four test cases (ROBER, POLLU, H2 autoignition, and H2–NH3 turbulent jet flame) demonstrated the effectiveness of physical constraints, BCT, and the balanced loss function.
@@
Figure 10: AE integrated DeepONet for stiff chemical kinetics. Reprinted from Reference 139, with permission.
@@
Some studies used NNs to learn chemical tabulations, which can be viewed as surrogate modeling of chemical kinetics as well. In References 270, 271, 272, NNs were used to learn the tabulation of thermochemistry following the paradigm of Equation (4.13). The hybrid flamelet/random data were generated for training and multiple NNs were used to predict each species with different magnitudes of concentration changes. The species conservation (Equation (3.1)) was enforced by normalizing the predicted species concentrations. Readshaw et al. 273 extended this framework by imposing constraints on element conservation, which was achieved by recalculating the mass fractions of selected residual species after the initial prediction. The results demonstrated that this conserved NN exhibited higher accuracy and physical consistency than the non-conserved NNs. Salunkhe et al. 274 developed the framework of ChemTab to jointly learn the progress variable generation (PVG) and manifold approximation (MA). For the PVG, the high dimensional species mass fractions are linearly embedded to the progress variables: $\mathbf{C_{pv}}=\mathbf{YW}$, while MA means the mapping between $\mathbf{C_{pv}}\oplus Z$ and the source terms is learned by a NN, where $Z$ is the mixture fraction. Therefore, the overall framework is similar to Equation (4.15). Inspired by the principal component analysis (PCA), some constraints were added in the loss function, including the orthogonality of $\mathbf{W}$ and $\mathbf{C_{pv}}\oplus Z$, and the unit-norm constraint on $\mathbf{W}$. It was demonstrated that removing any of these constraints will make the results worse. Salunkhe et al. 275 extended the ChemTab by adding another NN to learn the relationship between $\mathbf{C_{pv}}\oplus Z$ and lower dimensional source terms, which was computed using the same projection method and the same $\mathbf{W}$, as illustrated in Figure 11. The unit-norm constraint on $\mathbf{W}$ was replaced with the non-negativity constraint, which was achieved in a hard manner by truncating the negative values.
@@
Figure 11: ChemTab extended architecture and training procedure. Reprinted from Reference 275, with permission.
@@ 4.3 Neural ODEs: surrogate models and kinetics discovery
Neural ODE (NODE) was first proposed by Chen et al. 288 to develop a new family of NNs that parameterize the derivative of the hidden state using NNs, which is different from the traditional approach that defines discrete sequence of hidden layers, such as the ResNet and RNNs. Equation (4.14) is the basic expression of NODEs, while state variables are computed by integrating the Equation (4.14):
@@
Compared to discrete approaches, NODEs can better learn the continuous dynamics of a system, thus showing better reconstruction and extrapolation performance 288. Since NODE is proposed for learning the first-order dynamics of sequential systems, it is natural to apply it to chemical kinetic systems. Extensive studies have applied NODEs to the modeling of chemical kinetics for both building surrogate models and discovering kinetics, where many of them have informed/embedded physics to the NODEs.
@@ 4.3.1 Continuous surrogate models
As mentioned in Section 4.2, NODEs can be viewed as continuous surrogate models of chemical kinetics following the paradigm of Equation (4.14). Owoyele et al. 54 first applied NODEs to model chemical kinetics, termed as ChemNODE and shown in Figure 12. To improve the training stability, a progressive training strategy was employed, where the species were trained sequentially. For a homogeneous hydrogen-air reactor, ChemNODE showed accurate results across various operating conditions and yielded a speedup of 2.3x over the full chemical mechanism method. Kim et al. 289 used NODEs to learn stiff ODEs. The tests on the ROBER and POLLU problems demonstrated that the scaling of equations and loss functions was essential for NODEs to learn stiff ODEs.
@@
Figure 12: The ChemNODE approach. $\bm{\psi}$ indicates the thermochemical scalars. Reprinted from Reference 54, with permission.
@@
Although the name contains the "ODE", the vanilla NODE is indeed a data-driven method because no information of the ODEs is utilized. Recall that PINNs predict $\mathbf{y}$ by $\mathbf{y}=\mathrm{NN}(t)$ and differentiate it to get $\dot{\mathbf{y}}$, while NODEs, in contrast, predict $\dot{\mathbf{y}}$ by $\dot{\mathbf{y}}=\mathrm{NN}(\mathbf{y}(t))$ and integrate it to get $\mathbf{y}$. Therefore, if the ODEs can be explicitly added to the loss functions of NODEs, they can be viewed as PINNs in a broad sense, or an integral version of PINNs. Some studies have made the NODEs physics-informed in different ways. Sholokhov et al. 290 developed the physics-informed NODE (PI-NODE) to build reduced-order models (ROMs) for dynamical systems, where an AE was employed to perform model reduction and the NODE was used to learn the nonlinear dynamics in the latent space. Based on the ODEs in the physical space, the latent-space dynamics was derived and constitute a physics-informed loss, which was evaluated on a set of collocation points. The addition of the physics-informed loss improved the long-term performance, the robustness to data scarcity and noise, and the compressibility of the ROMs. Vijayarangan et al. 276 developed an AE-based ROM for stiff chemical kinetics using dynamics-informed training, where the NODE was used to learn the latent-space dynamics. The "dynamics-informed" means that the discovery of the latent space is informed by the latent dynamical evolution rule during the training procedure, which is different from the definition of "physics-informed" in a narrow sense. Two additional losses were added to ensure the mapping of the encoder and decoder was bijective. Results demonstrated that the proposed ROM could serve as an effective stiffness-removal framework. Kumar et al. 277, 278 developed the physics-constrained NODE (PC-NODE) to learn stiff chemical kinetics, where the conservations of each element were constrained by including corresponding loss terms. Here the element conservation was constrained between the predicted and true values, instead of the values at time $t$ and $t+\Delta t$ in the paradigm of discrete surrogate models. Compared to the purely data-driven NODE, results demonstrated that the PC-NODE framework improved not only the physical consistency but also the training efficiency. Later, Kumar et al. 279 combined the NODE with AE to develop the Phy-ChemNODE. The four loss terms can be considered as a combination of the bijective constraints of the dynamics-informed NODE 276 and the element conservation constraints of the PC-NODE 277, 278. Phy-ChemNODE was tested on a CH4-O2 combustion kinetics, where the 33 physical variables were reduced to 4 latent variables. The element conservation constraints were shown to help the model to obey the mass conservation better.
@@
Although the vanilla NODEs are purely data-driven, their interpretability can be enhanced by using the KANs as the network architecture. Koenig et al. 291 proposed the KAN-ODEs, combining the advantage of NODEs that do not require prior knowledge and the advantages of KANs that have increased interpretability and faster neural scaling laws compared to MLPs. The KAN-ODE was validated on learning Lotka-Volterra ODEs, modeling source terms of the Fisher-KPP PDE, and learning 1D PDEs. The scaling performance was shown superior to MLPs and the model performance could be further enhanced by sparsification and using symbolic regression. Subsequently, based on the KAN-ODEs and LeanKAN 292, Koenig et al. 280 proposed the ChemKAN for modeling and acceleration of combustion chemical kinetics. The physics was informed in two ways. First, the architecture strongly mimicked the chemical kinetics, as illustrated in Figure 13. A KAN was first used to represent the species production rates (core structure), then they were linearly combined to represent the temperature rate (thermodynamic superstructure) with another KAN to reflect the thermophysical parameter variation. This serial architecture offers the flexibility that one can only learn the species when temperature is not a variable. Second, an optional loss function for element conservation was introduced, which is similar to the PC-NODE and Phy-ChemNODE. The first test case was a biodiesel production modeling, where the thermodynamic superstructure and physical loss were "turned off". Compared with DeepONet, ChemKAN exhibited stronger scaling behaviour on test sets, less tendency toward overfitting, and stronger robustness to noise. In the second case of H2-air combustion, the ability of ChemKAN for computation acceleration was validated. Compared to MLP-based ChemNODE, ChemKAN showed similar speed-up ratios, but solved more species.
@@
Figure 13: The structure of ChemKAN. Reprinted from Reference 280, with permission.
@@
In addition to inform physics in the form of loss functions, some studies embed the information of stoichiometry and/or rate expressions into the NODEs to enhance the physical consistency. According to Equation (3.29), the rates of species and reactions are linearly related by the stoichiometric matrix. Thus, if the stoichiometric matrix is known, one can use NODEs to directly predict $\mathbf{RoP}$ and then transform it to $\dot{\bm{\upomega}}$, or vice versa. Based on the known information, Sorourifar et al. 117 developed two versions of physics-enhanced NODEs for chemical reaction systems: r-NODE and k-NODE. The r-NODE approximates $\mathbf{RoP}$ when only the stoichiometry is known and embedded, while the k-NODE approximates the kinetic constants when the mass-action parts of $\mathbf{RoP}$ are further known and embedded. Results suggest that a ground truth model that accurately reflects the data can be learned only when the embedded information accurately reflects the underlying mechanics. Fedorov et al. 118 developed the kinetics-constrained NODEs (KC-NODEs) for kinetic modeling, where the kinetic and thermodynamic knowledge were embedded into the NN architecture, including the stoichiometric matrix, the law of mass action, and the Arrhenius dependency of the rates on temperature. The KC-NODE was successfully applied to the kinetic modeling of the hydrogenation of CO2 to methane using numerical data and to higher hydrocarbons using real experimental data. Kircher et al. 119, 120 proposed the global reaction neural network (GRNN) to learn kinetics from reactor data, where the stoichiometry and thermodynamics were explicitly embedded as the network layers, as shown in Figure 14. The GRNN can be trained directly when source data are provided and trained in a NODE way when only reactor data are available, which is a 1D plug flow reactor model in the study. It was demonstrated that the embedded knowledge enhanced the model robustness to significant noise in the data.
@@
Figure 14: GRNN architecture with embedded stoichiometry and thermodynamics. Reprinted from Reference 119.

### v2

@@ 4.2 Surrogate models of chemical kinetics
In combustion simulations, the direct integration of the chemical reaction source terms is computationally expensive and is the main bottleneck for fast simulations of industrial combustion. Therefore, many methods have been proposed to accelerate the computation 44, 45, 116, 272, such as tabulation methods, reduced chemistry, and manifold methods. ML methods have also been adopted for computation acceleration, usually by constructing NN surrogate models of the chemical kinetics 44. The NNs are usually first trained using kinetic data, and then used to replace the integration of reaction source terms, so the overall computation efficiency can be improved. Based on the type of NN outputs, the NN surrogate models can be classified into two paradigms: discrete and continuous models. Specifically, the discrete models learn the mapping between the state variables at the current time and next time (i.e., following the AR paradigm), or the corresponding changes:
@@
Almeldein et al. 270 developed the physics-informed surrogate NNs for accelerating chemical kinetics calculations following Equation (4.12), where the conservation of four elements was added in the loss function. The mixture-of-experts architecture was adopted to allow multiple subnetworks to specialize in different regimes, such as pre-ignition low-temperature chemistry, high-temperature reactions, and post-combustion equilibriums. The 0D CH4-air combustion was tested, demonstrating the effectiveness of element conservation and the acceleration ability of the surrogate models on both CPUs and GPUs. Zhang et al. 271, 288 proposed the parametric physics-informed surrogate NNs for combustion reaction kinetics (CRK-PINNs) based on the paradigm of Equation (4.12). As illustrated in Figure 8, the kinetic ODEs and the conservations of elements, species, and energy were considered in the loss function to improve the physical completeness and low data dependence. A logarithmic normalization strategy was adopted to mitigate the multiscale of species and stiffness of the system. Based on the H2-air mechanism with 10 species and 21 reactions, the CRK-PINN was validated in three scenarios: solving forward and inverse problems, surrogate chemical kinetics, and accelerating simulations of combustion reacting flows. In the first scenario, CRK-PINN can solve the ignition process without data, reconstruct the process with sparse data, and infer intermediate species and temperature with the data of major species. In the second scenario, CRK-PINN can surrogate chemical kinetics in wide operating conditions and exhibited higher accuracy and lower data dependence than data-driven NNs for both time-continuous and time-advanced predictions. In the last scenario, CRK-PINN can replace the direct integration solver to predict the reaction source terms in the DNS of a 2D laminar Bunsen flame and a 3D turbulent jet flame, showing superior accuracy than data-driven NNs and great acceleration compared to the direct integration method.
Figure 8: CRK-PINN architecture. Reprinted from Reference 271, with permission.
In addition to the soft constraints in the form of loss functions, there are also studies that impose hard constraints on physical conservations. Zhang et al. 117 proposed a multiscale sampling method to enhance the accuracy and robustness of NNs to surrogate chemical kinetics, showing superior ability than the manifold, Monte Carlo, and GAN sampling. The Box-Cox transformation (BCT) was adopted to preprocess the data of species mass fractions:
@@
which is a good alternative to the logarithmic normalization 271. Although the paradigm of Equation (4.13) was adopted, only the species were directly predicted by the NNs, while the temperature and pressure were updated by transport fluxes, thus ensuring the mass and energy conservation. The algorithm has been validated on various CFD cases of combustion and been applied to DeepFlame 118, a deep learning empowered open-source platform for reacting flow simulations. Wang et al. 116 extended the method by imposing constraints on element conservations in both soft and hard ways. As illustrated in Figure 9, the ANN-soft penalized the element conservations by adding a loss term, while ANN-hard enforced the element conservations by adding a correction term, $\Delta\mathbf{Y}_{c}$, to the NN-predicted mass fractions. The correction term can be easily derived by the element conservations and the changes of species mass fractions, and was first computed in the Box-Cox transformed space to avoid the multiscale issue. One advantage of ANN-hard over ANN-soft is that there is no cumbersome adjustment of the loss weights. In terms of prediction accuracy and error of element conservations, ANN-hard exhibited superior performance than ANN-soft and they all outperformed the regular ANN on various flame cases. The computational overhead caused by the correction term was negligible, demonstrating the efficiency of ANN-hard. In fact, the correction method to ensure element conservations has been adopted in earlier studies 54, 55, where the NNs were trained to surrogate the kinetics of reduced chemistry, following the paradigm of Equation (4.15).s
@@
Figure 9: ANN-hard: schematic and comparisons with other approaches. Reprinted from Reference 116, with permission.
@@
In addition to the paradigm of function learning, some studies follow the paradigm of operator learning to learn combustion chemical kinetics, using neural operators such as the DeepONet 98 and FNO 99. Some data-driven DeepONets have been used to surrogate chemical kinetics 289, 142, 143, where the input of branch nets and the final output are similar to Equation (4.12), while the input of truck nets is related to the temporal coordinates or increments, enabling the models to learn kinetics with variable $\Delta t$, which is often fixed in classical surrogate models. To some extent, they can be viewed as a hybrid paradigm of AR and NAR learning. One insight that these works can provide is that learning kinetics in the latent space with reduced dimension can improve computational efficiency, which can be achieved by integrating a self-supervised autoencoder (AE) 142, 143, as shown in Figure 10. Kumar et al. 143 also accelerated the computation by selecting representative species for kinetics learning. The other scalars can be reconstructed by a pre-trained correlation NN. This strategy was also adopted in Reference 105, 106, where the species and element conservations were integrated into the loss functions of both correlation NN and DeepONet, thus constituting a physics-informed learning. The framework was tested for CH4 oxidation, demonstrating that the physical constraints can reduce the conservation errors without loss of prediction accuracy. Zanardi et al. 107 used the PI-DeepONets to learn the solution operator of non-equilibrium chemistry in hypersonic flows, where the coarse-grained (CG) kinetic ODEs were considered. The input of the branch and truck nets were the ICs and temporal coordinates, respectively. The framework was tested on the O2–O mixture, exhibiting two orders of magnitude in speedup compared to numerical integrators. Later, Zanardi et al. 108 extended the framework to be more physics-informed by many strategies, including a Boltzmann transformation layer to enforce local equilibrium distributions between states in the same group, the hierarchical architecture and transfer learning between different temporal scales, and the adaptive online pruning technique. Weng et al. 272 applied FNO to learn stiff chemical kinetics, following the paradigm of Equation (4.12). The loss of element and species conservations were considered and the BCT (Equation (4.16)) was also adopted. Based on the focal loss 290, a balanced loss function was proposed to address the unbalanced sampling issue in complex reacting flow problems. The results on four test cases (ROBER, POLLU, H2 autoignition, and H2–NH3 turbulent jet flame) demonstrated the effectiveness of physical constraints, BCT, and the balanced loss function.
@@
Figure 10: AE integrated DeepONet for stiff chemical kinetics. Reprinted from Reference 142, with permission.
@@
Some studies used NNs to learn chemical tabulations, which can be viewed as surrogate modeling of chemical kinetics as well. In References 273, 274, 275, NNs were used to learn the tabulation of thermochemistry following the paradigm of Equation (4.13). The hybrid flamelet/random data were generated for training and multiple NNs were used to predict each species with different magnitudes of concentration changes. The species conservation (Equation (3.1)) was enforced by normalizing the predicted species concentrations. Readshaw et al. 276 extended this framework by imposing constraints on element conservation, which was achieved by recalculating the mass fractions of selected residual species after the initial prediction. The results demonstrated that this conserved NN exhibited higher accuracy and physical consistency than the non-conserved NNs. Salunkhe et al. 277 developed the framework of ChemTab to jointly learn the progress variable generation (PVG) and manifold approximation (MA). For the PVG, the high dimensional species mass fractions are linearly embedded to the progress variables: $\mathbf{C_{pv}}=\mathbf{YW}$, while MA means the mapping between $\mathbf{C_{pv}}\oplus Z$ and the source terms is learned by a NN, where $Z$ is the mixture fraction. Therefore, the overall framework is similar to Equation (4.15). Inspired by the principal component analysis (PCA), some constraints were added in the loss function, including the orthogonality of $\mathbf{W}$ and $\mathbf{C_{pv}}\oplus Z$, and the unit-norm constraint on $\mathbf{W}$. It was demonstrated that removing any of these constraints will make the results worse. Salunkhe et al. 278 extended the ChemTab by adding another NN to learn the relationship between $\mathbf{C_{pv}}\oplus Z$ and lower dimensional source terms, which was computed using the same projection method and the same $\mathbf{W}$, as illustrated in Figure 11. The unit-norm constraint on $\mathbf{W}$ was replaced with the non-negativity constraint, which was achieved in a hard manner by truncating the negative values.
@@
Figure 11: ChemTab extended architecture and training procedure. Reprinted from Reference 278, with permission.
@@ 4.3 Neural ODEs: surrogate models and kinetics discovery
Neural ODE (NODE) was first proposed by Chen et al. 291 to develop a new family of NNs that parameterize the derivative of the hidden state using NNs, which is different from the traditional approach that defines discrete sequence of hidden layers, such as the ResNet and RNNs. Equation (4.14) is the basic expression of NODEs, while state variables are computed by integrating the Equation (4.14):
@@
Compared to discrete approaches, NODEs can better learn the continuous dynamics of a system, thus showing better reconstruction and extrapolation performance 291. Since NODE is proposed for learning the first-order dynamics of sequential systems, it is natural to apply it to chemical kinetic systems. Extensive studies have applied NODEs to the modeling of chemical kinetics for both building surrogate models and discovering kinetics, where many of them have informed/embedded physics to the NODEs.
@@ 4.3.1 Continuous surrogate models
As mentioned in Section 4.2, NODEs can be viewed as continuous surrogate models of chemical kinetics following the paradigm of Equation (4.14). Owoyele et al. 57 first applied NODEs to model chemical kinetics, termed as ChemNODE and shown in Figure 12. To improve the training stability, a progressive training strategy was employed, where the species were trained sequentially. For a homogeneous hydrogen-air reactor, ChemNODE showed accurate results across various operating conditions and yielded a speedup of 2.3x over the full chemical mechanism method. Kim et al. 292 used NODEs to learn stiff ODEs. The tests on the ROBER and POLLU problems demonstrated that the scaling of equations and loss functions was essential for NODEs to learn stiff ODEs.
@@
Figure 12: The ChemNODE approach. $\bm{\psi}$ indicates the thermochemical scalars. Reprinted from Reference 57, with permission.
@@
Although the name contains the "ODE", the vanilla NODE is indeed a data-driven method because no information of the ODEs is utilized. Recall that PINNs predict $\mathbf{y}$ by $\mathbf{y}=\mathrm{NN}(t)$ and differentiate it to get $\dot{\mathbf{y}}$, while NODEs, in contrast, predict $\dot{\mathbf{y}}$ by $\dot{\mathbf{y}}=\mathrm{NN}(\mathbf{y}(t))$ and integrate it to get $\mathbf{y}$. Therefore, if the ODEs can be explicitly added to the loss functions of NODEs, they can be viewed as PINNs in a broad sense, or an integral version of PINNs. Some studies have made the NODEs physics-informed in different ways. Sholokhov et al. 293 developed the physics-informed NODE (PI-NODE) to build reduced-order models (ROMs) for dynamical systems, where an AE was employed to perform model reduction and the NODE was used to learn the nonlinear dynamics in the latent space. Based on the ODEs in the physical space, the latent-space dynamics was derived and constitute a physics-informed loss, which was evaluated on a set of collocation points. The addition of the physics-informed loss improved the long-term performance, the robustness to data scarcity and noise, and the compressibility of the ROMs. Vijayarangan et al. 279 developed an AE-based ROM for stiff chemical kinetics using dynamics-informed training, where the NODE was used to learn the latent-space dynamics. The "dynamics-informed" means that the discovery of the latent space is informed by the latent dynamical evolution rule during the training procedure, which is different from the definition of "physics-informed" in a narrow sense. Two additional losses were added to ensure the mapping of the encoder and decoder was bijective. Results demonstrated that the proposed ROM could serve as an effective stiffness-removal framework. Kumar et al. 280, 281 developed the physics-constrained NODE (PC-NODE) to learn stiff chemical kinetics, where the conservations of each element were constrained by including corresponding loss terms. Here the element conservation was constrained between the predicted and true values, instead of the values at time $t$ and $t+\Delta t$ in the paradigm of discrete surrogate models. Compared to the purely data-driven NODE, results demonstrated that the PC-NODE framework improved not only the physical consistency but also the training efficiency. Later, Kumar et al. 282 combined the NODE with AE to develop the Phy-ChemNODE. The four loss terms can be considered as a combination of the bijective constraints of the dynamics-informed NODE 279 and the element conservation constraints of the PC-NODE 280, 281. Phy-ChemNODE was tested on a CH4-O2 combustion kinetics, where the 33 physical variables were reduced to 4 latent variables. The element conservation constraints were shown to help the model to obey the mass conservation better.
@@
Although the vanilla NODEs are purely data-driven, their interpretability can be enhanced by using the KANs as the network architecture. Koenig et al. 294 proposed the KAN-ODEs, combining the advantage of NODEs that do not require prior knowledge and the advantages of KANs that have increased interpretability and faster neural scaling laws compared to MLPs. The KAN-ODE was validated on learning Lotka-Volterra ODEs, modeling source terms of the Fisher-KPP PDE, and learning 1D PDEs. The scaling performance was shown superior to MLPs and the model performance could be further enhanced by sparsification and using symbolic regression. Subsequently, based on the KAN-ODEs and LeanKAN 295, Koenig et al. 283 proposed the ChemKAN for modeling and acceleration of combustion chemical kinetics. The physics was informed in two ways. First, the architecture strongly mimicked the chemical kinetics, as illustrated in Figure 13. A KAN was first used to represent the species production rates (core structure), then they were linearly combined to represent the temperature rate (thermodynamic superstructure) with another KAN to reflect the thermophysical parameter variation. This serial architecture offers the flexibility that one can only learn the species when temperature is not a variable. Second, an optional loss function for element conservation was introduced, which is similar to the PC-NODE and Phy-ChemNODE. The first test case was a biodiesel production modeling, where the thermodynamic superstructure and physical loss were "turned off". Compared with DeepONet, ChemKAN exhibited stronger scaling behaviour on test sets, less tendency toward overfitting, and stronger robustness to noise. In the second case of H2-air combustion, the ability of ChemKAN for computation acceleration was validated. Compared to MLP-based ChemNODE, ChemKAN showed similar speed-up ratios, but solved more species.
@@
Figure 13: The structure of ChemKAN. Reprinted from Reference 283, with permission.
@@
In addition to inform physics in the form of loss functions, some studies embed the information of stoichiometry and/or rate expressions into the NODEs to enhance the physical consistency. According to Equation (3.29), the rates of species and reactions are linearly related by the stoichiometric matrix. Thus, if the stoichiometric matrix is known, one can use NODEs to directly predict $\mathbf{RoP}$ and then transform it to $\dot{\bm{\upomega}}$, or vice versa. Based on the known information, Sorourifar et al. 120 developed two versions of physics-enhanced NODEs for chemical reaction systems: r-NODE and k-NODE. The r-NODE approximates $\mathbf{RoP}$ when only the stoichiometry is known and embedded, while the k-NODE approximates the kinetic constants when the mass-action parts of $\mathbf{RoP}$ are further known and embedded. Results suggest that a ground truth model that accurately reflects the data can be learned only when the embedded information accurately reflects the underlying mechanics. Fedorov et al. 121 developed the kinetics-constrained NODEs (KC-NODEs) for kinetic modeling, where the kinetic and thermodynamic knowledge were embedded into the NN architecture, including the stoichiometric matrix, the law of mass action, and the Arrhenius dependency of the rates on temperature. The KC-NODE was successfully applied to the kinetic modeling of the hydrogenation of CO2 to methane using numerical data and to higher hydrocarbons using real experimental data. Kircher et al. 122, 123 proposed the global reaction neural network (GRNN) to learn kinetics from reactor data, where the stoichiometry and thermodynamics were explicitly embedded as the network layers, as shown in Figure 14. The GRNN can be trained directly when source data are provided and trained in a NODE way when only reactor data are available, which is a 1D plug flow reactor model in the study. It was demonstrated that the embedded knowledge enhanced the model robustness to significant noise in the data.
@@
Figure 14: GRNN architecture with embedded stoichiometry and thermodynamics. Reprinted from Reference 122.